OpenStudy (anonymous):

How would I use this optimized equation: V= 4x^3-64x^2+240x to find a max/min? I'm having trouble finding c

OpenStudy (anonymous):

you looking for the max and min of V?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

calculus class?

OpenStudy (anonymous):

Yep. V is the volume of a box

OpenStudy (anonymous):

And there is no given interval or anything

OpenStudy (anonymous):

take the derivative, set it equal to zero, solve for \(x\)

OpenStudy (anonymous):

it is a cubic polynomial so it goes from \(-\infty\) to \(\infty\) there will be no global max or min, only local ones

OpenStudy (anonymous):

That's where I'm stuck, I got 12x^2-128x+240=0, but I'm having trouble with the rest

OpenStudy (anonymous):

get rid of the 12 first

OpenStudy (anonymous):

meaning factor it out?

OpenStudy (anonymous):

ok i guess 12 does not go in to 128 but 4 does

OpenStudy (anonymous):

yeah the roots are ugly, need to use the quadratic formula that, or cheat

OpenStudy (anonymous):

Okay. I divided both sides by 4 (is that okay? because I'm dividing 0 by 4, so I'll lose it forever, right?) and got 3x^2-32x+60=0

OpenStudy (anonymous):

OH I forgot about the quadratic formula

OpenStudy (anonymous):

no real choices here you can try to factor all day and you will not find the factors use the formula

OpenStudy (anonymous):

So I'd get a plus/minus something and then I'd plug it into the original equation, and the smaller number would be the min and the larger number would be the max?

OpenStudy (anonymous):

oh no you don't need to do that

OpenStudy (anonymous):

cubic polynomial, leading coefficient is positive, looks like this 

OpenStudy (anonymous):

the smaller of the two roots will be a local max, the larger will be a local min you know this before you start

OpenStudy (anonymous):

Oh okay. Does that work for all critical values, or just for cubic functions?

OpenStudy (anonymous):

no not for everything my guess is you have seen a cubic polynomial so you have a general idea of what they look like of course 4th degree polynomials look different etc

OpenStudy (anonymous):

Okay! In other cases though, I'd plug it into the original, etc. right?

OpenStudy (anonymous):

this like one of these "a box is made out of a flat sheet of cardboard..." problems?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

lol how'd i guess? usually they are cooked up to give nice answers, i guess not in this case

OpenStudy (anonymous):

did you start with something like \[V(x)=x(8-2x)(12-2x)\] ?

OpenStudy (anonymous):

No, it was V= x(12-2x)(20-2x)

OpenStudy (anonymous):

ok so one side was 12 and the other side 20 too bad this one gives nice ugly answers

OpenStudy (anonymous):

Yeah, it is! Anyways, thanks for the help!!

OpenStudy (anonymous):

yw

OpenStudy (amistre64):

was this part of a larger question?